OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -5x-10/-5.. the whole expression is over -5

OpenStudy (anonymous):

Tere are a number of approaches to solve this problem-- this is just one.\[\frac{-5x-10}{-5}=-\frac{1}{5}(-5x-10)=\frac{1}{5}*5x+\frac{1}{5}*10=x+2\]Note: after the distribution I dropped the negative signs because (-1)(-1)=+1

OpenStudy (anonymous):

thanks